Suppose that the cost function for a pickle cannery is given by TC= 9000 1000

Info I found from answers I got for other questions:

MC=q/500; AC=q/1000+9000/q; Firm size that minimizes average cost-->q=3000; Long Run Price-->P=AC=$6; Equilibrium quantity of pickles consumed=30,000; number of canaries in the industry=10; short-run supply-->q=500P

A) What is the quantity of pickles consumed in the short run?
Answer #1

The number of canneries which will lead to the optimum average cost is d(AC)/dq=1/1000-9000/q^2

Equating it to 0, we get q^2=9, so q=3

That means the numbers are given in thousands. (As, 3000=3x1000)

Since the number of canaries is 10, in the short run, everything will be consumed. Thus, the quantity will be 10x1000=10,000